# Towards Efficient PCSEL Design: A Data-Driven Approach for Design Insights

### v3

@@ Abstract
We present a data-driven design approach for photonic crystals to achieve high efficiency in photonic crystal surface-emitting lasers (PCSELs). By discretizing the photonic crystal structure into a grid, we enable the generation of arbitrary lattice designs. Multiple fully connected layers combined with a position embedding module extract essential features from the photonic crystal designs, while coupled-wave theory (CWT) is used to evaluate the efficiency (based on the ratio of surface-emitting to edge-emitting resonant) and quality factor Q. We introduce the Neural Networks (NNs) model to evaluate the structures, and to find a better performance design according to the evaluation result. The model achieves high prediction accuracy, with Pearson correlation coefficients of 0.780 for SEE and 0.887 for the log-transformed Q. Additionally, we perform Shapley value analysis to identify the most important Fourier coefficients, providing insights into the factors that impact the performance of PCSEL designs. Our work speeds up the design process and offers valuable guidance for optimizing high-performance PCSELs, supporting the development of fully photonic design automation (PDA).
@@ Introduction
Photonic crystals are structures with periodic dielectric materials that create photonic band structures, enabling the control of light propagation precisely. These structures have been widely used in various applications, including waveguides, filters, and lasers [1, 2, 3]. Photonic crystal surface-emitting laser (PCSEL), an important application of photonic crystals, has attracted significant interest due to their potential across a wide range of applications, including light detection and ranging (LIDAR) systems, laser processing, and optical communications [4, 5]. These devices offer the unique advantage of operating in a surface-emitting configuration, which provides several benefits over traditional semiconductor lasers, such as edge-emitting lasers (EELs) and vertical cavity surface-emitting lasers (VCSELs). These benefits include high output power, excellent beam quality, an ultra-small divergence angle, and single-mode operation [6, 7, 8]. Despite these advantages, traditional methods for designing photonic crystals often rely on trial-and-error or parameter sweep approaches [9, 10, 11, 12, 13]. These methods can be both time-consuming and computationally expensive, making it difficult to efficiently explore the vast design space of PCSELs.
Data-driven approaches have gained traction in the field of photonics, as they provide a more efficient alternative for design optimization [14, 15]. These methods leverage machine learning models to explore complex design spaces and identify optimal solutions without the need for exhaustive simulations. By applying these techniques, the design process can be significantly accelerated, and the computational burden reduced [16, 17]. However, most existing approaches focus on general optimization rather than providing insights into the specific features of photonic crystals that impact PCSEL, which limits their applicability to high-efficiency PCSEL design.
In this study, we propose a novel data-driven approach tailored specifically for the design of PCSELs. By discretizing the photonic crystal structure into a grid, our method facilitates the generation of a wide variety of lattice configurations, significantly expanding the range of possible designs. One of the key features of our approach is the use of Neural Networks (NNs) [18], a type of machine learning model well-suited for identifying intricate patterns and relationships within complex datasets. NNs excel at extracting essential features from the photonic crystal structures, such as Fourier coefficients, which serve as a compact representation of the structure’s key characteristics. By leveraging NNs, we are able to capture critical features, such as symmetry and destructive interactions, in double-lattice [10] or triple-lattice photonic crystals [19], that strongly influence laser performance.
@@
In addition to the performance prediction, we perform Shapley value analysis to assess the relative importance of each Fourier coefficient in determining the efficiency of PCSEL designs. Shapley values, derived from cooperative game theory [23], offer a quantitative measure of each input parameter’s impact on the model’s output. This analysis reveals the most influential features of the photonic crystal structure, providing valuable insights into the key factors that drive the performance of PCSELs.
By combining data-driven design methods, CWT, and Shapley value analysis, our approach not only speeds up the design process but also offers a deeper understanding of the structural features that impact PCSEL efficiency. This method provides a powerful tool for optimizing PCSELs and offers the potential for greater automation in photonic device design. In the following sections, we describe the implementation of our approach, present the results from model training and testing, and discuss the insights gained from Shapley value analysis.
@@ Methods
Photonic crystal simulation
To design arbitrary photonic crystals, we first define a random filling factor ($FF$), which determines the ratio of high and low dielectric constant materials within the photonic crystal and is expressed as $FF=S_{\mathrm{low\ dielectric}}/S_{\mathrm{high\ dielectric}}$. The unit cell of the photonic crystal is discretized into an $n\times n$ grid (for a square lattice). For each grid cell, a value $z=f(x,y)$ is generated, with $f(x,y)$ defining the spatial distribution. This value is compared with a threshold based on the filling factor (FF). If the generated value exceeds the threshold, the grid cell is assigned to a high dielectric constant material (e.g., GaAs). Otherwise, it is assigned to a low dielectric constant material (e.g., air).
For the numerical analysis of the photonic crystal, we apply Coupled-Wave Theory (CWT) to analyze a square lattice structure. In this approach, we consider the propagation of four basic waves along the $x$- and $y$-directions relative to the photonic crystal. The coupling between these waves can be described using the coupling matrix C, which can be expressed in the following form:
@@
Thus, the we can describe the four basic waves in the photonic crystal as:
@@
Applying the boundary conditions $R_{x}(0,y)=S_{x}(L,y)=R_{y}(x,0)=S_{y}(x,L)=0$ for a finite-size $L\times L$ square-lattice photonic crystal, we can solve Eq. (5) to obtain the spatial distribution of four basic waves. We can then calculate the optical power of the surface-emitting and edge-emitting, and derive the efficiency of the PCSEL. The optical power of all stimulated emission, surface-emitting and emitting can be expressed as:
@@
To evaluate the efficiency of the designed PCSEL, we consider the ratio of surface-emitting power to the stimulated emission power as surface-emitting efficiency ($SEE$). Since we are designing a surface-emitting laser, a higher ratio of $SEE=P_{\mathrm{surface}}/P_{\mathrm{stim}}$ indicates higher efficiency of the laser. Specifically, the efficiency is enhanced when the surface-emitting power is maximized relative to the total stimulated emission, which leads to greater power extraction through the surface of the device. This ratio can be used as a key performance indicator for optimizing the photonic crystal structure for high-efficiency laser operation.
In addition to surface-emitting efficiency, the value of the quality factor $Q$ is also an important metric for evaluating the performance of PCSELs. The quality factor is defined as the ratio of the energy stored in the cavity to the energy lost per cycle, and is given by $Q=\frac{2\pi/a}{\alpha_{\mathrm{r}}}$, where $a$ is the lattice constant of the photonic crystal. A high quality factor indicates low cavity loss and high energy storage, which are essential for low-threshold lasing. Therefore, the quality factor is also a key result that we need to extract from the model to ensure the laser operates at its optimal performance.To handle the large variation in the value of $Q$ (ranging from $10^{3}$ to $10^{5}$), we trained the model to predict $\log{Q}$ instead of $Q$ directly. This transformation reduces the dynamic range of the target values, thereby stabilizing the training process and improving convergence. After prediction, the final $Q$ value is obtained by taking the exponential of the model output. This approach ensures that the model captures the underlying pattern more effectively while minimizing sensitivity to extreme variations in $Q$.
@@
We developed a fully connected neural network (shown in Figure 1), to predict $SEE$ and $\log{Q}$ of PCSEL designs. The model begins by 1024 input grid, which represents the real and imaginary parts of the Fourier coefficients ($\xi^{X}_{m,n}$, where $X\in\{R,I\}$) of the discretized photonic crystal structure. A learnable positional embedding [24] ($PE^{R}_{m,n}$ and $PE^{I}_{m,n}$) is then added to this flattened input to help capture spatial information. We set $PE^{R}_{m,n}=PE^{I}_{m,n}$ to ensure that the real and imaginary parts of the Fourier coefficients have the same positional embedding. For simplicity and to reduce redundancy, we only retain one of the pairings of $\xi^{X}_{m,n}$ and $\xi^{X}_{-m,-n}$ in our analysis. This is because these coefficients are complex conjugates of each other and represent the same input. By removing the conjugate pair, we avoid duplicating the information, ensuring more efficient processing without loss of relevant data.
@@
This architecture is designed to effectively learn the mapping between the grid-based representation of the photonic crystal and its performance, enabling accurate prediction of the laser’s performance.
@@
The dataset for training and testing was generated using a Python-based simulation pipeline. In this process, random parameters are first selected to define the characteristics of three holes in the photonic crystal unit cell. For each hole, random center positions, standard deviations, and orientation angles are generated. A two-dimensional Gaussian function is then used to produce a spatial distribution for each hole, and these distributions are summed to form a composite sample array. The final resulting dielectric array is reshaped and averaged into a $32\times 32$ grid.
The simulation function is subsequently called with the generated dielectric array and a solver instance. This function performs optical simulations based on coupled-wave theory (CWT) and finite element method (FEM) analysis, and returns key performance metrics such as the $SEE$ and quality factor $Q$, and other relevant parameters. We adopt the epitaxial structure outlined in Table 1, with the Bragg wavelength fixed at 980 nm. To ensure consistency with the finite-size PCSELs, the photonic crystal’s lateral dimension $L$ is chosen to be as close as possible to 200 $\mu m$, while maintaining an integer multiple of the crystal period. The photonic crystal unit cells are discretized using a $17\times 17$ grid to capture the structural details accurately. The results are then written to a NPZ file (see supplementary information), which serves as the dataset for further model training and testing. This iterative process continues until an external stop signal is detected, thereby building a comprehensive dataset of photonic crystal designs. Finally, our dataset consists of 200k simulation results, ensuring a diverse and representative set of samples for training robust machine learning models.
@@
The epitaxial structure used in this work, as shown in Table 1, serves as an example for the design of the PCSEL. The choice of epitaxial layers mainly impacts the Green’s function and the confinement factor of the photonic crystal ($\Gamma_{\mathrm{PhC}}$) within the framework of CWT. However, since the primary focus of this work is on the design of the photonic crystal structure itself, the specific epitaxial layers chosen do not significantly affect the overall performance. Furthermore, the methodology proposed here can be easily adapted to other epitaxial structures, making the design approach flexible for various PCSELs epitaxial structures.
@@
We trained a fully connected neural network to predict key performance metrics of photonic crystal surface-emitting laser (PCSEL) designs based on the generated dataset. The dataset, was split into 190k samples for training and 10k samples for testing. The model was trained using the Adam optimizer and training process was run for up to 100 epochs, with early stopping based on the validation loss to prevent overfitting (shown in Figure 2 (a) and (d)). Train epoch finally stopped at about 45 epochs. We evaluated the model’s performance on the test set using the Pearson correlation coefficient as the primary metric. Pearson’s correlation coefficient is used to evaluate the linear relationship between predicted and actual values, providing insight into the model’s predictive accuracy. The results are summarized in Table 2, which shows a high correlation between the predicted and actually simulated efficiency values. In Figure 2 (b) and (e), we present the test results for $SEE$ and $\log{Q}$, respectively. These results demonstrate that the model effectively captures the key performance characteristics of PCSEL designs, with relatively high prediction accuracy. The error distribution for the training results of $SEE$ and $\log{Q}$ is shown in Figure 2 (c) and (f), respectively. About 80% of the predictions fall within a range of ±$\pm 0.14$ for $SEE$ and ±$\pm 0.15$ for $\log{Q}$, indicating that the model produces accurate estimates with minimal deviation from the simulated values.
@@
As shown in Table 2 these results demonstrate that the neural network model effectively captures the key performance characteristics of PCSEL designs. The Pearson correlation coefficients which means the linear relationship between predicted and actual values, are 0.780 for $SEE$ and 0.887 for $\log{Q}$. These high correlation values indicate that the model has strong predictive capability. The relatively low RMSE (Root of Mean Square Error) values further suggest that the model produces accurate estimates with a minimal deviation from the simulated values. We also calculated the coefficient of determination ($R^{2}$) to evaluate the proportion of variance in the data that is explained by the model. A $R^{2}$ value closer to 1 indicates that the model explains a larger portion of the variance in the data, confirming its predictive power. In our results, the $R^{2}$ values of 0.595 for $SEE$ and 0.784 for $\log{Q}$ confirm that the model explains a substantial portion of the variance in the data. These findings highlight the potential of data-driven approaches for accelerating the optimization of PCSEL designs, reducing reliance on computationally expensive simulations while maintaining high prediction accuracy.
Global optimization of photonic crystal design
After establishing the predictive capabilities of our trained neural network models for surface-emitting efficiency ($SEE$) and quality factor ($Q$), we leverage these models to search for a photonic crystal design that minimizes $\left|SEE_{Target}-SEE\right|$ while ensuring $Q>10^{4}$ to maintain a low operational threshold.
@@
We employ a global optimization algorithm, specifically the DIRECT [25, 26] (DIviding RECTangles) method, to efficiently explore the parameter space. The optimization variables include the positions, sizes, and orientations of the holes in the photonic crystal unit cell, as well as the filling factor. By utilizing the trained models for $SEE$ and $Q$, the optimization process is significantly accelerated, as each evaluation bypasses the need for relatively computationally expensive coupled-wave theory (CWT) simulations. To validate the optimal design obtained via the model-driven optimization, we perform full physical simulations using the CWT and FEM solvers. A comparison between predicted and simulated results further confirms the reliability of the neural network models in guiding the optimization process. The optimized structure is summarized in Table 3. The final validation results demonstrate strong agreement with our model predictions. The entire optimization process involved over 16,000 function evaluations, with a total runtime of approximately only 2.2 minutes. This significantly reduced the computational burden compared to direct CWT simulations. This demonstrates the effectiveness of combining machine learning-based surrogate models with global optimization algorithms for rapid and accurate photonic crystal design.
@@
This approach demonstrates the potential of machine learning models not only in predicting photonic crystal performance but also in guiding global optimization for high-efficiency laser design.
@@
We employed Shapley values, a game-theoretic approach for interpreting machine learning models [23], to evaluate the influence of each input on the neural network model’s predictions. Since our input features are represented as the real and imaginary parts of the Fourier series coefficients of the photonic crystal structure, we computed Shapley values for each component separately to quantify their contribution to the predicted $SEE$ and $Q$. The Shapley value analysis was implemented using the SHAP library [27]. This approach helps us understand which Fourier coefficients most significantly affect the surface-emitting efficiency and quality factor, providing insight into the design of PCSELs.
To gain the relative impacts of different input features in determining the performance of PCSEL designs, we conducted a Shapley value analysis (on 500 randomly selected samples from the test dataset). The results of this analysis are shown in Figure 3, which illustrates the contribution of each Fourier coefficient to the predicted $SEE$ and $\log{Q}$. The Shapley values provide a quantitative measure of the impact of each input parameter on the model’s output, revealing the most influential features in the photonic crystal structure.
@@
The $\xi^{R}_{0,0}$, which corresponds to the average dielectric constant of the photonic crystal layer, also has a notable influence on both $SEE$ and $\log{Q}$. $\xi^{R}_{0,0}$ shows a generally positive correlation with both $SEE$ and $\log{Q}$. This suggests that higher values of $\xi^{R}_{0,0}$ lead to improved performance in both surface-emitting efficiency and the quality factor. We speculate that this is because $\xi^{R}_{0,0}$ directly influences the confinement factor ($\Gamma_{\mathrm{PhC}}$) within the photonic crystal, which in turn affects both the efficiency of surface-emitting power extraction and the quality factor. A higher $\xi^{R}_{0,0}$ means a better confinement of light within the structure, leading to enhanced optical performance, including higher $SEE$ and $\log{Q}$ values.

### v4

# Towards Efficient PCSEL Design: A Fully AI-driven Approach

@@ Abstract
We present an fully AI-driven design framework for photonic crystals (PhCs), engineered to achieve high efficiency in photonic crystal surface-emitting lasers (PCSELs). By discretizing the PhC structure into a grid, where the edges of the holes are represented by the cross-sections of two-dimensional Gaussian surfaces, we achieve high-degree-of-freedom and fabrication-friendly hole design. Coupled-wave theory (CWT) generates a dataset by evaluating surface-emitting efficiency ($SEE$) and quality factor ($Q$) of PhC designs, while a multi-layered neural network (NN) learns and extracts essential features from these designs. Finally, black-box optimization (BBO) is employed to fine-tune the photonic crystal structure, enabling a fully AI-driven design process. The model achieves high prediction accuracy, with Pearson correlation coefficients of 0.780 for $SEE$ and 0.887 for the log-transformed $Q$. Additionally, we perform Shapley value analysis to identify the most important Fourier coefficients, providing insights into the factors that impact the performance of PCSEL designs. Our work accelerates the design process by over 1,000,000 times compared to traditional FDTD simulations, reducing parameter optimization from two weeks to just one second. Our work speeds up the design process and enables efficient optimization of high-performance PCSELs, driving the development of fully photonic design automation (PDA).
@@ Introduction
Photonic crystals (PhCs) are structures with periodic dielectric materials that create photonic band structures, enabling the control of light propagation precisely. These structures have been widely used in various applications, including waveguides, filters, and lasers [1, 2, 3]. Photonic crystal surface-emitting laser (PCSEL), an important application of PhCs, has attracted significant interest due to their potential across a wide range of applications, including light detection and ranging (LIDAR) systems, laser processing, and optical communications [4, 5]. These devices offer the unique advantage of operating in a surface-emitting configuration, which provides several benefits over traditional semiconductor lasers, such as edge-emitting lasers (EELs) and vertical cavity surface-emitting lasers (VCSELs). These benefits include high output power, excellent beam quality, an ultra-small divergence angle, and single-mode operation [6, 7, 8]. Despite these advantages, traditional methods for designing PhCs often rely on trial-and-error or parameter sweep approaches [9, 10, 11, 12, 13]. These methods can be both time-consuming and computationally expensive, making it difficult to efficiently explore the vast design space of PCSELs.
AI-driven approaches have gained traction in the field of photonics, as they provide a more efficient alternative for design optimization [14, 15, 16]. These methods leverage machine learning models to explore complex design spaces and identify optimal solutions without the need for exhaustive simulations. By applying these techniques, the design process can be significantly accelerated, and the computational burden reduced [17, 18]. However, most existing approaches focus on general optimization rather than providing insights into the specific features of PhCs that impact PCSEL, which limits their applicability to high-efficiency PCSEL design.
In this study, we propose a fully AI-driven approach tailored specifically for the design of PCSELs. By discretizing the PhC structure into grids, our method facilitates the generation of a wide variety of lattice configurations, significantly expanding the range of possible designs. One of the key features of our approach is the use of Neural Networks (NNs) [19], a type of machine learning model well-suited for identifying intricate patterns and relationships within complex datasets. NNs excel at extracting essential features from the PhC structures, such as Fourier coefficients, which serve as a compact representation of the structure’s key characteristics. By leveraging NNs, we are able to capture critical features, such as symmetry and destructive interactions, in double-lattice [10] or triple-lattice PhCs [20], that strongly influence laser performance.
@@
In addition to the performance prediction, we perform Shapley value analysis to assess the relative importance of each Fourier coefficient in determining the efficiency of PCSEL designs. Shapley values, derived from cooperative game theory [24], offer a quantitative measure of each input parameter’s impact on the model’s output. This analysis reveals the most influential features of the PhC structure, providing valuable insights into the key factors that drive the performance of PCSELs.
By combining AI-driven design methods, CWT, and Shapley value analysis, our approach not only speeds up the design process but also offers a deeper understanding of the structural features that impact PCSEL efficiency. This method provides a powerful tool for optimizing PCSELs and offers the potential for greater automation in photonic device design. In the following sections, we describe the implementation of our approach, present the results from model training and testing, and discuss the insights gained from Shapley value analysis.
@@ Methods
PCSEL simulation
To design arbitrary PhCs, we first define a random filling factor ($FF$), which determines the ratio of high and low dielectric constant materials within the PhC and is expressed as $FF=S_{\mathrm{low\ dielectric}}/\left(S_{\mathrm{low\ dielectric}}+S_{\mathrm{high\ dielectric}}\right)$. The unit cell of the PhC is discretized into an $n\times n$ grid (for a square lattice). For each grid cell, a value $z=G(x,y)$ is generated using the two-dimensional Gaussian surface cross-section fitting method (2D-GCF, Eq. 15). This value is compared with a threshold $T_{0}$ based on the filling factor (FF). If $z$ larger than $T_{0}$, the grid cell is assigned to a high dielectric constant material (e.g., GaAs). Otherwise, it is assigned to a low dielectric constant material (e.g., air).
For the numerical analysis of the PCSEL, we apply Coupled-Wave Theory (CWT) to analyze a square lattice structure. In this approach, we consider the propagation of four basic waves along the $x$- and $y$-directions relative to the PhC. The coupling between these waves can be described using the coupling matrix C, which can be expressed in the following form:
@@
Thus, the we can describe the four basic waves in the PhC as:
@@
Applying the boundary conditions $R_{x}(0,y)=S_{x}(L,y)=R_{y}(x,0)=S_{y}(x,L)=0$ for a finite-size $L\times L$ square-lattice PhC, we can solve Eq. (5) to obtain the spatial distribution of four basic waves. We can then calculate the optical power of the surface-emitting and edge-emitting, and derive the efficiency of the PCSEL. The optical power of all stimulated emission, surface-emitting and emitting can be expressed as:
@@
To evaluate the efficiency of the designed PCSEL, we consider the ratio of surface-emitting power to the stimulated emission power as surface-emitting efficiency ($SEE$). Since we are designing a surface-emitting laser, a higher ratio of $SEE=P_{\mathrm{surface}}/P_{\mathrm{stim}}$ indicates higher efficiency of the laser. Specifically, the efficiency is enhanced when the surface-emitting power is maximized relative to the total stimulated emission, which leads to greater power extraction through the surface of the device. This ratio can be used as a key performance indicator for optimizing the PhC structure for high-efficiency laser operation.
In addition to surface-emitting efficiency, the value of the quality factor $Q$ is also an important metric for evaluating the performance of PCSELs. The quality factor is defined as the ratio of the energy stored in the cavity to the energy lost per cycle, and is given by $Q=\frac{2\pi/a}{\alpha_{\mathrm{r}}}$, where $a$ is the lattice constant of the PhC. A high quality factor indicates low cavity loss and high energy storage, which are essential for low-threshold lasing. Therefore, the quality factor is also a key result that we need to extract from the model to ensure the laser operates at its optimal performance.To handle the large variation in the value of $Q$ (ranging from $10^{3}$ to $10^{5}$), we trained the model to predict $\log{Q}$ instead of $Q$ directly. This transformation reduces the dynamic range of the target values, thereby stabilizing the training process and improving convergence. After prediction, the final $Q$ value is obtained by taking the exponential of the model output. This approach ensures that the model captures the underlying pattern more effectively while minimizing sensitivity to extreme variations in $Q$.
@@
We developed a fully connected neural network (shown in Figure 1), to predict $SEE$ and $\log{Q}$ of PCSEL designs. The model begins by 1024 input grid, which represents the real and imaginary parts of the Fourier coefficients ($\xi^{X}_{m,n}$, where $X\in\{R,I\}$) of the discretized PhC structure. A learnable positional embedding [25] ($PE^{R}_{m,n}$ and $PE^{I}_{m,n}$) is then added to this flattened input to help capture spatial information. We set $PE^{R}_{m,n}=PE^{I}_{m,n}$ to ensure that the real and imaginary parts of the Fourier coefficients have the same positional embedding. For simplicity and to reduce redundancy, we only retain one of the pairings of $\xi^{X}_{m,n}$ and $\xi^{X}_{-m,-n}$ in our analysis. This is because these coefficients are complex conjugates of each other and represent the same input. By removing the conjugate pair, we avoid duplicating the information, ensuring more efficient processing without loss of relevant data.
@@
This architecture is designed to effectively learn the mapping between the grid-based representation of the PhC and its performance, enabling accurate prediction of the laser’s performance.
@@
The dataset for training and testing was generated using the two-dimensional Gaussian surface cross-section fitting method (2D-GCF). In this approach, random parameters are assigned to define the characteristics of multiple two-dimensional Gaussian distributions. Specifically, for each Gaussian distribution, random values are generated for the amplitude $A$, center positions $(x_{0},y_{0})$, standard deviations $(\sigma_{x},\sigma_{y})$, and orientation angles $(\theta)$. These individual Gaussian distributions are then summed to form a composite function $G(x,y)$. Finally, the contour where $G(x,y)$ equals a specific threshold value $T_{0}$ is extracted to define the edges of the holes in the PhC structure. The resulting clear hole contours are pixelated into a $32\times 32$ grid map (dielectric array) for further processing. Due to the inherent smoothness of Gaussian functions, the designed hole contours are guaranteed to be smooth, which provides significant convenience for subsequent fabrication processes. Moreover, this randomization and summation process ensures a diverse and comprehensive dataset, capturing a wide range of possible configurations for robust model training and evaluation. In this study, the authors fixed the number of Gaussian distributions to three.
$g_{i}(x,y)=A_{i}\cdot\exp\left(-\frac{(x-x_{0,i})^{2}}{2\sigma_{x,i}^{2}}-\frac{(y-y_{0,i})^{2}}{2\sigma_{y,i}^{2}}\right),i\in\mathbb{Z}^{+}$ (14)
$G(x,y)=\sum_{i=1}^{n}g_{i}(x,y)$ (15)
$G(x,y)=T_{0}$ (16)
The simulation function is subsequently called with the generated dielectric array and a solver instance. This function performs optical simulations based on coupled-wave theory (CWT) and finite element method (FEM) analysis, and returns key performance metrics such as the $SEE$ and quality factor $Q$, and other relevant parameters. We adopt the epitaxial structure outlined in Table 1, with the Bragg wavelength fixed at 980 nm. To ensure consistency with the finite-size PCSELs, the PhC’s lateral dimension $L$ is chosen to be as close as possible to 200 $\mu m$, while maintaining an integer multiple of the crystal period. The finite-size PhC are discretized to a $17\times 17$ grid for FEM calculation. The results are then written to a NPZ file (see supplementary information), which serves as the dataset for further model training and testing. This iterative process continues until an external stop signal is detected, thereby building a comprehensive dataset of PhC designs. Finally, our dataset consists of 200k simulation results, ensuring a diverse and representative set of samples for training robust machine learning models.
@@
The epitaxial structure used in this work, as shown in Table 1, serves as an example for the design of the PCSEL. The choice of epitaxial layers mainly impacts the Green’s function and the confinement factor of the photonic crystal ($\Gamma_{\mathrm{PhC}}$) within the framework of CWT. However, since the primary focus of this work is on the design of the PhC structure itself, the specific epitaxial layers chosen do not significantly affect the overall performance. Furthermore, the methodology proposed here can be easily adapted to other epitaxial structures, making the design approach flexible for various PCSELs epitaxial structures.
@@
We trained a fully connected neural network to predict key performance metrics of PCSEL designs based on the generated dataset. The dataset was split into 190k samples for training and 10k samples for testing. The model was trained using the Adam optimizer and training process was run for up to 100 epochs, with early stopping based on the validation loss to prevent overfitting (shown in Figure 2 (a) and (d)). Train epoch finally stopped at about 45 epochs. We evaluated the model’s performance on the test set using the Pearson correlation coefficient as the primary metric. Pearson’s correlation coefficient is used to evaluate the linear relationship between predicted and actual values, providing insight into the model’s predictive accuracy. The results are summarized in Table 2, which shows a high correlation between the predicted and actually simulated efficiency values. In Figure 2 (b) and (e), we present the test results for $SEE$ and $\log{Q}$, respectively. These results demonstrate that the model effectively captures the key performance characteristics of PCSEL designs, with relatively high prediction accuracy. The error distribution for the training results of $SEE$ and $\log{Q}$ is shown in Figure 2 (c) and (f), respectively. About 80% of the predictions fall within a range of ±$\pm 0.14$ for $SEE$ and ±$\pm 0.15$ for $\log{Q}$, indicating that the model produces accurate estimates with minimal deviation from the simulated values.
@@
As shown in Table 2 these results demonstrate that the neural network model effectively captures the key performance characteristics of PCSEL designs. The Pearson correlation coefficients which means the linear relationship between predicted and actual values, are 0.780 for $SEE$ and 0.887 for $\log{Q}$. These high correlation values indicate that the model has strong predictive capability. The relatively low RMSE (Root of Mean Square Error) values further suggest that the model produces accurate estimates with a minimal deviation from the simulated values. We also calculated the coefficient of determination ($R^{2}$) to evaluate the proportion of variance in the data that is explained by the model. A $R^{2}$ value closer to 1 indicates that the model explains a larger portion of the variance in the data, confirming its predictive power. In our results, the $R^{2}$ values of 0.595 for $SEE$ and 0.784 for $\log{Q}$ confirm that the model explains a substantial portion of the variance in the data. These findings highlight the potential of fully AI-driven approaches for accelerating the optimization of PCSEL designs, reducing reliance on computationally expensive simulations while maintaining high prediction accuracy.
Global optimization of PhC design
After establishing the predictive capabilities of our trained neural network models for $SEE$ and $Q$, we leverage these models to search for a PhC design that minimizes $\left|SEE_{Target}-SEE\right|$ while ensuring $Q>10^{4}$ to maintain a low operational threshold.
@@
We employ a global black-box optimization (BBO) algorithm, specifically the DIRECT [26, 27] (DIviding RECTangles) method, to efficiently explore the parameter space. The optimization variables include the positions, sizes, and orientations of the holes in the PhC unit cell, as well as the filling factor. By utilizing the trained models for $SEE$ and $Q$, the optimization process is significantly accelerated, as each evaluation bypasses the need for relatively computationally expensive coupled-wave theory (CWT) simulations. To validate the optimal design obtained via the model-driven optimization, we perform full physical simulations using the CWT and FEM solvers. A comparison between predicted and simulated results further confirms the reliability of the neural network models in guiding the optimization process. The optimized structure is summarized in Table 3. The final validation results demonstrate strong agreement with our model predictions. The entire optimization process involved over 16,000 function evaluations, with a total runtime of approximately only 2.2 minutes. This significantly accelerates the design process by over 1,000,000 times compared to traditional FDTD simulations, which usually cost several hours per finite PCSEL simulation on standard workstation computers. This demonstrates the effectiveness of combining machine learning-based models with BBO algorithm for rapid and accurate PhC design.
@@
This approach demonstrates the potential of machine learning models not only in predicting PhC performance but also in guiding global optimization for high-efficiency laser design.
@@
We employed Shapley values, a game-theoretic approach for interpreting machine learning models [24], to evaluate the influence of each input on the neural network model’s predictions. Since our input features are represented as the real and imaginary parts of the Fourier series coefficients of the PhC structure, we computed Shapley values for each component separately to quantify their contribution to the predicted $SEE$ and $Q$. The Shapley value analysis was implemented using the SHAP library [28]. This approach helps us understand which Fourier coefficients most significantly affect the surface-emitting efficiency and quality factor, providing insight into the design of PCSELs.
To gain the relative impacts of different input features in determining the performance of PCSEL designs, we conducted a Shapley value analysis (on 500 randomly selected samples from the test dataset). The results of this analysis are shown in Figure 3, which illustrates the contribution of each Fourier coefficient to the predicted $SEE$ and $\log{Q}$. The Shapley values provide a quantitative measure of the impact of each input parameter on the model’s output, revealing the most influential features in the PhC structure.
@@
The $\xi^{R}_{0,0}$, which corresponds to the average dielectric constant of the PhC layer, also has a notable influence on both $SEE$ and $\log{Q}$. $\xi^{R}_{0,0}$ shows a generally positive correlation with both $SEE$ and $\log{Q}$. This suggests that higher values of $\xi^{R}_{0,0}$ lead to improved performance in both surface-emitting efficiency and the quality factor. We speculate that this is because $\xi^{R}_{0,0}$ directly influences the confinement factor ($\Gamma_{\mathrm{PhC}}$) within the PhC, which in turn affects both the efficiency of surface-emitting power extraction and the quality factor. A higher $\xi^{R}_{0,0}$ means a better confinement of light within the structure, leading to enhanced optical performance, including higher $SEE$ and $\log{Q}$ values.